# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 2.3.3 Encoder-decoder models
Evolving further, BART (Bidirectional and Auto-Regressive Transformers) was introduced by Lewis et al. from Facebook AI in 2019.102 BART combined the context learning strengths of the bidirectional BERT, and the autoregressive capabilities of models like GPT, which excel at generating coherent text. BART was thus a hybrid seq2seq model, consisting of a BERT-like bidirectional encoder and a GPT-like autoregressive decoder. This is nearly the same architecture as Vaswani et al. 69; the differences are in the pretraining. BART was pretrained using a task that corrupted text by, for example, deleting tokens, and shuffling sentences. It then learned to reconstruct the original text with left-to-right autoregressive decoding as in GPT models.
@@ 2.3.4 Multi-task and multi-modal models
In previous sections, we discussed LLMs that take natural language text as input and then they output either a learned representation or another text sequence. These models traditionally perform tasks like translation, summarization, and classification. Raffel et al. 103 developed the Text-to-Text Transfer Transformer (T5) to demonstrate that various tasks can be reframed as text-to-text tasks, allowing the same model architecture and training procedure to be used with the same set of weights across different tasks. This approach has inspired the development of multi-task models that can adapt to different tasks at inference time. For example, Flan-T5104 used instruction fine-tuning with chain-of-thought prompts, enabling it to generalize to unseen tasks, such as generating rationales before answering. More complex approaches have been proposed for robust multi-task models.105, 106, 107, 108
Additionally, LLMs have been extended to understand different input modalities, despite initially receiving only text. For instance, Fuyu109 uses linear projection to adapt image representations into the token space of an LLM, enabling a decoder-only model to write figure captions. Scaling up, next-GPT110 was developed as an “any-to-any” model, processing multiple modalities—text, audio, image, and video—through modality-specific encoders. The encoded representation is projected into a decoder-only token space, and the LLM’s output is processed by a domain-specific diffusion model to generate each modality’s output. Multitask or multimodel methods are further described below as these methods start to connect LLMs with autonomous agents.
@@ 3.1 Molecular Representations, Datasets, and Benchmarks
Given the diversity of data in chemistry-based machine learning, and the many different ways that a compound can be described, from molecular structure to property description, there are multiple ways a molecule can be represented, underlining this heterogeneity.111, 112, 113, 114 Some common forms include molecular graphs115, 116, 117, 3D point clouds118, 119, 120, 121, and quantitative feature descriptors.122, 123, 124, 125, 112 In this review, we will focus on string-based representations of molecules, given the interest in language models. Among the known string representations, we can cite IUPAC names, SMILES,41 DeepSMILES,126 SELFIES,127 and InChI,128 as recently reviewed by Das et al. 129.
@@
Figure 2: Number of training tokens (on log scale) available from various chemical sources compared with typical LLM training runs. The numbers are drawn from ZINC130, PubChem131, Touvron et al. 132, ChEMBL133, and Kinney et al. 134
@@
Regarding datasets, there are two types of data used for training LLMs: training data and evaluation data. Training data should be grounded on real molecules to build the correct “prior belief” of a molecule definition during the LLM training process. Similar care is taken when natural language training data used in GPT-4, for instance, is built from real sentences or code, not random combinations of words that might be grammatically correct. Figure 2 shows the number of tokens on common chemistry datasets in comparison with the LLaMA2 training data, extracted from References [130, 131, 132, 133, 134]. According to this data, the largest chemical training corpus, composed of hypothetical chemical structures, composes billions of tokens against a few trillion in the LLaMA 2 training data. Interestingly, Segler et al. 135 demonstrated that even using the Reaxys dataset, a very small human-curated set of chemical reactions, is enough to enable state-of-the-art retrosynthesis results. Therefore, the lack of data is not the only problem. Instead, the lack of good-quality data seems to be the pivotal factor holding back the development of better scientific LLMs. Figure 2 illustrates the number of tokens on ZINC, a dataset with hypothetical chemical structures130
@@
On the other hand, benchmark data evaluates the success of these models. Numerous datasets, curated by the scientific community, are available for this purpose136, 137. Among them, MoleculeNet55, first published in 2017, is the most commonly used labeled dataset for chemistry. However, MoleculeNet has several issues: it is small, contains errors and inconsistencies, and lacks relevance to real-world chemistry problems138, 139, 140, 141. Pat Walters, a leader in ML for drug discovery, stated, “I think the best way to make progress on applications of machine learning to drug discovery is to fund a large public effort that will generate high-quality data and make this data available to the community”142.
@@
Walters makes several constructive critiques such as, for example, how the QM7, QM8, and QM9 datasets, intended for predicting quantum properties from 3D structures, are often misused with predictions based incorrectly on 1D SMILES strings, which inadequately represent 3D molecular conformations. Walters also makes suggestions for more relevant benchmarks, and also datasets with more valid entries. He suggests the Absorption, Distribution, Metabolism, and Excretion (ADME) data curated by Fang et al. 143, and the Therapeutic Data Commons (TDC)144, 145 and TDC-2146. These datasets contain measurements of real compounds, making them grounded in reality. Moreover, ADME is crucial for determining a drug candidate’s success, while therapeutic results in diverse modalities align with metrics used in drug development.
@@
Here, we hypothesize that the lack of easily accessible, high-quality data in the correct format for training foundational chemical language models is a major bottleneck to the development of the highly desired “super-human” AI-powered digital chemist. In addition, we argue that extracting data from scientific chemistry papers might be an interesting commitment to generating data of high quality, grounded to the truth, and on a large scale.
{forest}
@@ 3.2 Property Prediction and Encoder-only mol-LLMs
Table 1: Encoder-only scientific LLMs. aa:“Model Size” is reported as the number of parameters. $b$: The authors report they not used as many encoder layers as it was used in the original BERT paper. But the total number of parameters was not reported.
@@
CatBERTa164 355M OpenCatalyst2020 (OC20) RoBERTa Property prediction 2023.11
SELFormer176 $\sim$86M $\sim$2M compounds from ChEMBL Property prediction RoBERTa 2023.05
MaterialsBERT184 110M 2.4M material science abstracts + 750 annotated abstract for NER BERT NER and property extraction 2023.04
SolvBERT185 b 1M SMILES of solute-solvent pairs from CombiSolv-QM and LogS from Boobier et al. 186 BERT Property prediction 2023.01
MolFormer187 $b$ PubChem and ZINC BERT Property prediction 2022.12
ChemBERTa-2177 5M - 46M 77M SMILES from PubChem RoBERTa Property prediction 2022.09
ScholarBERT165 340M, 770M Public.Resource.Org, Inc BERT Property prediction 2022.05
MatSciBERT160 110M $\sim$150K material science paper downloaded from Elsevier BERT NER and text classification 2022.05
MatBERT168 110M Abstracts from solid state articles and abstracts and methods from gold nanoparticle articles BERT NER 2022.04
PubMedBERT161 110M 14M abstracts from PubMed BERT NER, QA, and document classification 2021.10
Mol-BERT179 110M $\sim$4B SMILES from ZINC15 and ChEMBL27 BERT Property prediction 2021.09
ChemBERT169 110M $\sim$200k extracted using ChemDataExtractor BERT NER 2021.06
MolBERT181 $\sim$85M ChemBench BERT Property prediction 2020.11
ChemBERTa42 10M SMILES from PubChem RoBERTa Property prediction 2020.10
BioMegatron170 345M, 800M, 1.2B Wikipedia, CC-Stories, Real-News, and OpenWebtext Megatron-LM NER and QA 2020-10
BioBERT171 110M PubMed and PMC BERT NER and QA 2020.02
SciBERT162 110M 1.14M papers from Semantic Scholar BERT NER and sentence classification 2019.09
SMILES-BERT188 $b$ $\sim$18M SMILES from ZINC BERT Property prediction 2019.09
BlueBERT163 110M PubMed and MIMIC-III BERT NER, and document classification 2019.06
ClinicalBERT166 110M MIMIC-III BERT Pacient readmission probability 2019.04
Encoder-only transformer architectures solely comprise an encoder. Chemists can leverage the simpler encoder-only architecture for tasks where the primary goal is classification, efficient exploration of chemical space, and property prediction. Since encoder-only architectures are mostly applied to property prediction, we describe here the relative importance of this principal chemical challenge. Sultan et al. 189 also discussed the high importance of this task.
@@ 3.2.1 Property Prediction
The universal value of chemistry lies in identifying and understanding the properties of compounds to optimize their practical applications. In the pharmaceutical industry, therapeutic molecules interact with the body in profound ways.190, 191, 192 Understanding these interactions and modifying molecular structures to enhance those therapeutic benefits can lead to significant medical advancements.193 Similarly, in polymer science, material properties depend on chemical structure, polymer chain length, and packing,194 and a protein’s function similarly depends on its structure and folding. Historically, chemists have identified new molecules from natural products195 and screened them against potential targets196 to test their properties for diseases. Once a natural product shows potential, chemists synthesize scaled-up quantities for further testing or derivatization,197, 198, 199 a costly and labor-intensive process.200, 201 Traditionally, chemists have used their expertise to hypothesize the properties of new molecules derived from those natural products, hence aiming for the best investment of synthesis time and labor. To support the chemical industry in more accurate property prediction, computational chemistry has evolved.202 Quantum theoretical calculations predict certain properties reliably, while force-field-based Molecular Dynamics (MD)203 predict packing and crystal structures of large molecular ensembles, though both require substantial computational resources.204, 205, 206, 207, 208 Property prediction can now be enhanced through machine learning tools,209, 124, 210, 211 and more recent advancements in LLMs lead to effective property prediction without the extensive computational demands of quantum mechanics and MD calculations. Combined with human insight, AI can revolutionize material development, enabling the synthesis of new materials with a high likelihood of possessing desired properties for specific applications.
@@ 3.2.2 Encoder-only Mol-LLMs
Encoder-only models are exemplified by the BERT architecture, which is commonly applied in natural language sentiment analysis to extract deeper patterns from prose.212 The human chemist has been taught to look at a 2D image of a molecular structure and to recognize its chemical properties or classify the compound. Therefore, encoder-only models would ideally convert SMILES strings, empty of inherent chemical essence, into a vector representation, or latent space, reflecting those chemical properties. This vector representation can then be used directly for various downstream tasks.
@@
Schwaller et al. 213 used a BERT model to more accurately classify complex synthesis reactions by generating reaction fingerprints from raw SMILES strings, without the need to separate reactants from reagents in the input data, thereby simplifying data preparation. The BERT model achieved higher accuracy (98.2%) compared to the encoder-decoder model (95.2%) for classifying reactions. Accurate classification aids in understanding reaction mechanisms, vital for reaction design, optimization, and retrosynthesis. Toniato et al. 214 also used a BERT architecture to classify reaction types for downstream retrosynthesis tasks that enable manufacture of any molecular target. Further examples of BERT use include unsupervised reaction atom-to-atom mapping 215, 216. These chemical classifications would accelerate research and development in organic synthesis, described further below.
@@
Due to limited labeled data for molecular property prediction, Wang et al. proposed a semi-supervised SMILES-BERT model, pretrained on a large unlabeled dataset with a Masked SMILES Recovery task.182 The model was then fine-tuned for various molecular property prediction tasks, outperforming state-of-the-art methods in 2019 on three chosen datasets varying in size and property. This marked a shift from using BERT for reaction classification towards property prediction and drug discovery.
@@
Zhang et al. 217 also tackled the issue of limited property-labeled data and the lack of correlation between any two datasets labeled for different properties, hindering generalizability. They introduced multitask learning BERT (MTL-BERT), which used large-scale pretraining and multitask learning with unlabeled SMILES strings from ChEMBL.133 This approach mined contextual information and extracted key patterns from complex SMILES strings, improving model interpretability. The model was fine-tuned for relevant downstream tasks, achieving better performance than state-of-the-art methods in 2022 on 60 molecular datasets from ADMETlab218 and MoleculeNet.55
@@
In 2021, Li and Jiang179 introduced Mol-BERT, pretrained on four million unlabeled drug SMILES from the ZINC15219 and ChEMBL27133 databases to capture molecular substructure information for property prediction. Their work leveraged the underutilized potential of large unlabeled datasets like ZINC. Mol-BERT consisted of three components: PretrainingExtractor, Pretraining Mol-BERT, and Fine-Tuning Mol-BERT. It treated Morgan fingerprint fragments as “words” and compounds as “sentences,” using RDKit and the Morgan algorithm for canonicalization and substructure identification. This approach generated comprehensive molecular fingerprints from SMILES strings, used in a Masked Language Model (MLM) task for pretraining. Mol-BERT was fine-tuned on labeled samples, providing outputs as binary values or continuous scores for classification or regression, and it outperformed existing sequence and graph-based methods by at least 2% in ROC-AUC scores on Tox21, SIDER, and ClinTox datasets.55
@@
Ross et al. 178 introduced MoLFormer, a large-scale unsupervised BERT model for more accurate and faster molecular property predictions than Density Functional Theory calculations or wet-lab experiments. They trained MoLFormer with rotary positional embeddings on SMILES sequences of 1.1 billion unlabeled molecules from PubChem131 and ZINC219. Rotary positional encoding captures token positions more effectively than traditional methods, 69 improving modeling of sequence relationships. MoLFormer outperformed state-of-the-art GNNs on several classification and regression tasks from ten MoleculeNet55 datasets, while performing competitively on two others. It effectively learned spatial relationships between atoms, predicting various molecular properties, including quantum-chemical properties.
@@
With ChemBERTa, Chithrananda et al. 42 explored the impact of pretraining dataset size, tokenization strategy, and the use of SMILES or SELFIES, distinguishing their work from other BERT studies. They used HuggingFace’s RoBERTa transformer,220 and referenced a DeepChem55 tutorial for accessibility. Their results showed improved performance on downstream tasks (BBBP, ClinTox, HIV, Tox21 from MoleculeNet55) as the pretraining dataset size increased from 100K to 10M. Although ChemBERTa did not surpass state-of-the-art GNN-based baselines like Chemprop (which used 2048-bit Morgan Fingerprints from rdkit),221 the authors suggested that with expansion to larger datasets they would eventually beat those baselines. The authors compared Byte-Pair Encoder (BPE) with a custom SmilesTokenizer and its regex developed by Schwaller et al.222 while exploring tokenization strategies. They found the SmilesTokenizer slightly outperformed BPE, suggesting more relevant sub-word tokenization is beneficial. No difference was found between SMILES and SELFIES, but the paper highlighted how attention heads in transformers could be visualized with BertViz,223 showing certain neurons selective for functional groups. This study underscored the importance of appropriate benchmarking and addresses the carbon footprint of AI in molecular property prediction.
@@
In ChemBERTa-2, Ahmad et al. 177 aimed to create a foundational model applicable across various tasks. They addressed criticisms of biased data in LLMs by training on 77M samples and adding a Multi-Task Regression component to the pretraining. ChemBERTa-2 matched state-of-the-art architectures on MoleculeNet.55 As with ChemBERTa, the work was valuable because of additional exploration, in this case into how pretraining improvements affected certain downstream tasks more than others, depending on the type of fine-tuning task, the structural features of the molecules in the fine-tuning task data set, or the size of that fine-tuning dataset. The result was that pretraining the encoder-only model is important, but gains could be made by considering the chemical application itself, and the associated fine-tuning dataset.

### v2

@@ 2.3.3 Encoder-decoder models
Evolving further, BART (Bidirectional and Auto-Regressive Transformers) was introduced by Lewis et al. from Facebook AI in 2019.109 BART combined the context learning strengths of the bidirectional BERT, and the autoregressive capabilities of models like GPT, which excel at generating coherent text. BART was thus a hybrid seq2seq model, consisting of a BERT-like bidirectional encoder and a GPT-like autoregressive decoder. This is nearly the same architecture as Vaswani et al. 71; the differences are in the pretraining. BART was pretrained using a task that corrupted text by, for example, deleting tokens, and shuffling sentences. It then learned to reconstruct the original text with left-to-right autoregressive decoding as in GPT models.
@@ 2.3.4 Multi-task and multi-modal models
In previous sections, we discussed LLMs that take natural language text as input and then they output either a learned representation or another text sequence. These models traditionally perform tasks like translation, summarization, and classification. Raffel et al. 110 developed the Text-to-Text Transfer Transformer (T5) to demonstrate that various tasks can be reframed as text-to-text tasks, allowing the same model architecture and training procedure to be used with the same set of weights across different tasks. This approach has inspired the development of multi-task models that can adapt to different tasks at inference time. For example, Flan-T5111 used instruction fine-tuning with chain-of-thought prompts, enabling it to generalize to unseen tasks, such as generating rationales before answering. More complex approaches have been proposed for robust multi-task models.112, 113, 114, 115
Additionally, LLMs have been extended to understand different input modalities, despite initially receiving only text. For instance, Fuyu116 uses linear projection to adapt image representations into the token space of an LLM, enabling a decoder-only model to write figure captions. Scaling up, next-GPT117 was developed as an “any-to-any” model, processing multiple modalities—text, audio, image, and video—through modality-specific encoders. The encoded representation is projected into a decoder-only token space, and the LLM’s output is processed by a domain-specific diffusion model to generate each modality’s output. Multitask or multimodel methods are further described below as these methods start to connect LLMs with autonomous agents.
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [SciLLMs, fill=blue, root [Text2Text, task [Question-answering, application [BioMedLM118, BioMistral119, BiMediX120, EpilepsyLLM121, CheXagent122, BioMedGPT-LM123, Darwin124, PMC-LLaMA125, Galactica126, BioGPT127, nacho0128, etc., study] ] ] [Text2Mol, task [Conditional de-novo generation, application [BioT5+129, Darwin124, Text+Chem T5130, MolT5131, etc., study] ] ] [Text2Number, task [Document classification, application [MatSciBERT132, Galactica126, PubMedBERT133, SciBERT134, BlueBERT135, etc, study] ] [Property prediction, application [EpilepsyLLM121, CatBERTa136, ScholarBERT137, ClinicalBERT138, etc, study] ] [Named entity recognition, application [nacho0128, MaterialsBert139, MatBERT140, ChemBERT141, PubMedBERT133, BioMegatron142, BioBERT143, SciBERT134, BlueBERT135, etc, study] ] ] [Mol2Text, task [Molecule captioning, application [BioT5+129, BioT5144, Text+Chem T5130, Galactica126, MolT5131, etc., study] ] ] [Mol2Mol, task [Molecule tuning, application [nacho0128, Regression Transformer145, ChemFormer146, etc., study] ] [Synthesis prediction, application [ReactionT5147, Galactica126, ChemFormer146, T5Chem148, etc., study] ] [Retrosynthesis, application [Text+Chem T5130, T5Chem148, etc., study] ] ] [Mol2Number, task [Property prediction, application [nacho0128, Regression Transformer145, CatBERTa136, Darwin124, SELFormer149, Galactica126, ChemBERTa-2150, MolFormer151, Mol-BERT152, MTL-BERT153, ChemBERTa44, MolBERT154, SMILES-BERT155, SMILES transformer156etc, study] ] ] ]
@@ 3.1 Molecular Representations, Datasets, and Benchmarks
Given the diversity of data in chemistry-based machine learning, and the many different ways that a compound can be described, from molecular structure to property description, there are multiple ways a molecule can be represented, underlining this heterogeneity.157, 158, 159, 160 Some common forms include molecular graphs161, 162, 163, 3D point clouds164, 165, 166, 167, and quantitative feature descriptors.168, 169, 170, 171, 158 In this review, we will focus on string-based representations of molecules, given the interest in language models. Among the known string representations, we can cite IUPAC names, SMILES,43 DeepSMILES,172 SELFIES,173 and InChI,174 as recently reviewed by Das et al. 175.
@@
Figure 3: Number of training tokens (on log scale) available from various chemical sources compared with typical LLM training runs. The numbers are drawn from ZINC176, PubChem177, Touvron et al. 178, ChEMBL179, and Kinney et al. 180
@@
Regarding datasets, there are two types of data used for training LLMs: training data and evaluation data. Training data should be grounded on real molecules to build the correct “prior belief” of a molecule definition during the LLM training process. Similar care is taken when natural language training data used in GPT-4, for instance, is built from real sentences or code; it is important not to use random combinations of words, even if they are grammatically correct. Figure 3 shows the number of tokens on common chemistry datasets in comparison with the number of tokens used to train LLaMA2, based on literature information.176, 177, 178, 179, 180 With this in mind, we note the largest chemical training corpus, which largely comprises hypothetical chemical structures, amounts to billions of tokens, almost two orders of magnitude fewer than the trillions of tokens used to train LLaMA2. When not including the large number of hypothetical structures found on ZINC176, (Figure 3), that number of tokens associated with verifiably synthesized compounds is over five orders of magnitude lower than the LLaMA2 training data. Nonetheless, Segler et al. 181 demonstrated that even using the Reaxys dataset, a very small human-curated set of chemical reactions, is enough to enable state-of-the-art retrosynthesis results. Therefore, the lack of data is not the only problem. Instead, the lack of good-quality data may be the pivotal factor holding back the development of better scientific LLMs.
@@
On the other hand, benchmark data evaluates the success of these models. Numerous datasets, curated by the scientific community, are available for this purpose182, 183. Among them, MoleculeNet57, first published in 2017, is the most commonly used labeled dataset for chemistry. However, MoleculeNet has several issues: it is small, contains errors and inconsistencies, and lacks relevance to real-world chemistry problems184, 146, 185, 128. Pat Walters, a leader in ML for drug discovery, stated, “I think the best way to make progress on applications of machine learning to drug discovery is to fund a large public effort that will generate high-quality data and make this data available to the community”186.
@@
Walters makes several constructive critiques such as, for example, how the QM7, QM8, and QM9 datasets, intended for predicting quantum properties from 3D structures, are often misused with predictions based incorrectly on 1D SMILES strings, which inadequately represent 3D molecular conformations. Walters also makes suggestions for more relevant benchmarks and also datasets with more valid entries. He suggests the Absorption, Distribution, Metabolism, and Excretion (ADME) data curated by Fang et al. 187, and the Therapeutic Data Commons (TDC)188, 189 and TDC-2190. These datasets contain measurements of real compounds, making them grounded in reality. Moreover, ADME is crucial for determining a drug candidate’s success, while therapeutic results in diverse modalities align with metrics used in drug development.
@@
Here, we hypothesize that the lack of easily accessible, high-quality data in the correct format for training foundational chemical language models is a major bottleneck to the development of the highly desired “super-human” AI-powered digital chemist. A more optimistic view is presented by Rich and Birnbaum 191. They argue that we do not need to wait for the creation of new benchmarks. Instead, the messy, available public data can be used to curate benchmarks carefully curated for specific real-world approximation. In addition, we argue that extracting data from scientific chemistry papers might be an interesting commitment to generating data of high quality, grounded to the truth, and on a large scale. Recently, a benchmark following these ideas — LAB-Bench192 — was created for evaluating LLMs’ performance in biology.
@@ 3.2 Property Prediction and Encoder-only mol-LLMs
Table 1: Encoder-only scientific LLMs. $a$:“Model Size” is reported as the number of parameters. $b$: The authors report they not used as many encoder layers as it was used in the original BERT paper. But the total number of parameters was not reported.
@@
CatBERTa136 355M OpenCatalyst2020 (OC20) RoBERTa Property prediction 2023.11
SELFormer149 $\sim$86M $\sim$2M compounds from ChEMBL RoBERTa Property prediction 2023.05
MaterialsBERT193 110M 2.4M material science abstracts + 750 annotated abstract for NER BERT NER and property extraction 2023.04
SolvBERT194 b 1M SMILES of solute-solvent pairs from CombiSolv-QM and LogS from Boobier et al. 195 BERT Property prediction 2023.01
MolFormer196 $b$ PubChem and ZINC BERT Property prediction 2022.12
ChemBERTa-2150 5M - 46M 77M SMILES from PubChem RoBERTa Property prediction 2022.09
ScholarBERT137 340M, 770M Public.Resource.Org, Inc BERT Property prediction 2022.05
BatteryBERT197 $\sim$ 110M $\sim$ 400k papers from RSC, Elsevier and Springer BERT Document classification 2022.05
MatSciBERT132 110M $\sim$150K material science paper downloaded from Elsevier BERT NER and text classification 2022.05
MatBERT140 110M Abstracts from solid state articles and abstracts and methods from gold nanoparticle articles BERT NER 2022.04
PubMedBERT133 110M 14M abstracts from PubMed BERT NER, QA, and document classification 2021.10
Mol-BERT152 110M $\sim$4B SMILES from ZINC15 and ChEMBL27 BERT Property prediction 2021.09
ChemBERT141 110M $\sim$200k extracted using ChemDataExtractor BERT NER 2021.06
MolBERT154 $\sim$85M ChemBench BERT Property prediction 2020.11
ChemBERTa44 10M SMILES from PubChem RoBERTa Property prediction 2020.10
BioMegatron142 345M, 800M, 1.2B Wikipedia, CC-Stories, Real-News, and OpenWebtext Megatron-LM NER and QA 2020-10
BioBERT143 110M PubMed and PMC BERT NER and QA 2020.02
Molecule Attention Transformer198 $b$ ZINC15 Encoder with GCN features Property prediction 2020.02
SciBERT134 110M 1.14M papers from Semantic Scholar BERT NER and sentence classification 2019.09
SMILES-BERT199 $b$ $\sim$18M SMILES from ZINC BERT Property prediction 2019.09
BlueBERT135 110M PubMed and MIMIC-III BERT NER, and document classification 2019.06
ClinicalBERT138 110M MIMIC-III BERT Patient readmission probability 2019.04
Encoder-only transformer architectures solely comprise an encoder. Chemists can leverage the simpler encoder-only architecture for tasks where the primary goal is classification, efficient exploration of chemical space, and property prediction. Since encoder-only architectures are mostly applied to property prediction, we describe here the relative importance of this principal chemical challenge. Sultan et al. 200 also discussed the high importance of this task, the knowledge obtained in the last years, and the remaining challenges regarding molecular property prediction using LLMs.
@@ 3.2.1 Property Prediction
The universal value of chemistry lies in identifying and understanding the properties of compounds to optimize their practical applications. In the pharmaceutical industry, therapeutic molecules interact with the body in profound ways.201, 202, 203 Understanding these interactions and modifying molecular structures to enhance those therapeutic benefits can lead to significant medical advancements.204 Similarly, in polymer science, material properties depend on chemical structure, polymer chain length, and packing,205 and a protein’s function similarly depends on its structure and folding. Historically, chemists have identified new molecules from natural products206 and screened them against potential targets207 to test their properties for diseases. Once a natural product shows potential, chemists synthesize scaled-up quantities for further testing or derivatization,208, 209, 210 a costly and labor-intensive process.211, 212 Traditionally, chemists have used their expertise to hypothesize the properties of new molecules derived from those natural products, hence aiming for the best investment of synthesis time and labor. To support the chemical industry in more accurate property prediction, computational chemistry has evolved.213 Quantum theoretical calculations predict certain properties reliably, while force-field-based Molecular Dynamics (MD)214 predict packing and crystal structures of large molecular ensembles, though both require substantial computational resources.215, 216, 217, 218, 219 Property prediction can now be enhanced through machine learning tools,220, 170, 221, 222 and more recent advancements in LLMs lead to effective property prediction without the extensive computational demands of quantum mechanics and MD calculations. Combined with human insight, AI can revolutionize material development, enabling the synthesis of new materials with a high likelihood of possessing desired properties for specific applications.
@@ 3.2.2 Encoder-only Mol-LLMs
Encoder-only models are exemplified by the BERT architecture, which is commonly applied in natural language sentiment analysis to extract deeper patterns from prose.223 The human chemist has been taught to look at a 2D image of a molecular structure and to recognize its chemical properties or classify the compound. Therefore, encoder-only models would ideally convert SMILES strings, empty of inherent chemical essence, into a vector representation, or latent space, reflecting those chemical properties. This vector representation can then be used directly for various downstream tasks.
@@
Schwaller et al. 224 used a BERT model to more accurately classify complex synthesis reactions by generating reaction fingerprints from raw SMILES strings, without the need to separate reactants from reagents in the input data, thereby simplifying data preparation. The BERT model achieved higher accuracy (98.2%) compared to the encoder-decoder model (95.2%) for classifying reactions. Accurate classification aids in understanding reaction mechanisms, vital for reaction design, optimization, and retrosynthesis. Toniato et al. 225 also used a BERT architecture to classify reaction types for downstream retrosynthesis tasks that enable manufacture of any molecular target. Further examples of BERT use include unsupervised reaction atom-to-atom mapping 226, 227. These chemical classifications would accelerate research and development in organic synthesis, described further below.
@@
Due to limited labeled data for molecular property prediction, Wang et al. proposed a semi-supervised SMILES-BERT model, pretrained on a large unlabeled dataset with a Masked SMILES Recovery task.155 The model was then fine-tuned for various molecular property prediction tasks, outperforming state-of-the-art methods in 2019 on three chosen datasets varying in size and property. This marked a shift from using BERT for reaction classification towards property prediction and drug discovery. Maziarka et al. 198 also claimed state-of-the-art performance in property prediction after self-supervised pretraining in their Molecule Attention Transformer (MAT), which adapted BERT to chemical molecules by augmenting the self-attention with inter-atomic distances and molecular graph structure.
@@
Zhang et al. 228 also tackled the issue of limited property-labeled data and the lack of correlation between any two datasets labeled for different properties, hindering generalizability. They introduced multitask learning BERT (MTL-BERT), which used large-scale pretraining and multitask learning with unlabeled SMILES strings from ChEMBL.179 This approach mined contextual information and extracted key patterns from complex SMILES strings, improving model interpretability. The model was fine-tuned for relevant downstream tasks, achieving better performance than state-of-the-art methods in 2022 on 60 molecular datasets from ADMETlab229 and MoleculeNet.57
@@
In 2021, Li and Jiang152 introduced Mol-BERT, pretrained on four million unlabeled drug SMILES from the ZINC15230 and ChEMBL27179 databases to capture molecular substructure information for property prediction. Their work leveraged the underutilized potential of large unlabeled datasets like ZINC. Mol-BERT consisted of three components: PretrainingExtractor, Pretraining Mol-BERT, and Fine-Tuning Mol-BERT. It treated Morgan fingerprint fragments as “words” and compounds as “sentences,” using RDKit and the Morgan algorithm for canonicalization and substructure identification. This approach generated comprehensive molecular fingerprints from SMILES strings, used in a Masked Language Model (MLM) task for pretraining. Mol-BERT was fine-tuned on labeled samples, providing outputs as binary values or continuous scores for classification or regression, and it outperformed existing sequence and graph-based methods by at least 2% in ROC-AUC scores on Tox21, SIDER, and ClinTox datasets.57
@@
Ross et al. 151 introduced MoLFormer, a large-scale unsupervised BERT model for more accurate and faster molecular property predictions than Density Functional Theory calculations or wet-lab experiments. They trained MoLFormer with rotary positional embeddings on SMILES sequences of 1.1 billion unlabeled molecules from PubChem177 and ZINC230. Rotary positional encoding captures token positions more effectively than traditional methods, 71 improving modeling of sequence relationships. MoLFormer outperformed state-of-the-art GNNs on several classification and regression tasks from ten MoleculeNet57 datasets, while performing competitively on two others. It effectively learned spatial relationships between atoms, predicting various molecular properties, including quantum-chemical properties. Additionally, the authors stated how MoLFormer represents an efficient and environment-friendly use of computational resources, reducing GPU usage in training by a factor of 60 (16 GPUs instead of 1000).
@@
With ChemBERTa, Chithrananda et al. 44 explored the impact of pretraining dataset size, tokenization strategy, and the use of SMILES or SELFIES, distinguishing their work from other BERT studies. They used HuggingFace’s RoBERTa transformer,231 and referenced a DeepChem57 tutorial for accessibility. Their results showed improved performance on downstream tasks (BBBP, ClinTox, HIV, Tox21 from MoleculeNet57) as the pretraining dataset size increased from 100K to 10M. Although ChemBERTa did not surpass state-of-the-art GNN-based baselines like Chemprop (which used 2048-bit Morgan Fingerprints from rdkit),232 the authors suggested that with expansion to larger datasets they would eventually beat those baselines. The authors compared Byte-Pair Encoder (BPE) with a custom SmilesTokenizer and its regex developed by 233 while exploring tokenization strategies. They found the SmilesTokenizer slightly outperformed BPE, suggesting more relevant sub-word tokenization is beneficial. No difference was found between SMILES and SELFIES, but the paper highlighted how attention heads in transformers could be visualized with BertViz,234 showing certain neurons selective for functional groups. This study underscored the importance of appropriate benchmarking and addresses the carbon footprint of AI in molecular property prediction.
@@
In ChemBERTa-2, Ahmad et al. 150 aimed to create a foundational model applicable across various tasks. They addressed a criticism that LLMs were not so generalizable because the training data was biased or non-representative. They addressed this criticism by training on 77M samples and adding a Multi-Task Regression component to the pretraining. ChemBERTa-2 matched state-of-the-art architectures on MoleculeNet.57 As with ChemBERTa, the work was valuable because of additional exploration, in this case into how pretraining improvements affected certain downstream tasks more than others, depending on the type of fine-tuning task, the structural features of the molecules in the fine-tuning task data set, or the size of that fine-tuning dataset. The result was that pretraining the encoder-only model is important, but gains could be made by considering the chemical application itself, and the associated fine-tuning dataset.